OpenStudy (anonymous):

If f(x) = {(3,5), (2,4),(1,7)} g(x) = sqrt of x-3 k(x) = x^2+5 determine: 1) composition - (g*k)(7) 2) 1/f(x) 3) (kg)(x)

OpenStudy (anonymous):

\[g\circ k(7)=g(k(7))=g(54)=\sqrt{21}\]

OpenStudy (anonymous):

\[\frac{1}{f(x)}:=\{(3,\frac{1}{5}),(2,\frac{1}{4}), (1,\frac{1}{7})\}\]

OpenStudy (anonymous):

and if the last one is multiplication just stick them next to each other with parentheses and a times sign between them

OpenStudy (anonymous):

so will it be (\[(x ^{2}+5)(\sqrt{x-3})\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks....=) Can you help me with another problem

OpenStudy (anonymous):

i made a typo on the first one!

OpenStudy (anonymous):

\[g\circ k(7)=g(k(7))=g(54)=\sqrt{51}\]

OpenStudy (anonymous):

thanks for fixing it...can i ask u another problem?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

or just post either way

OpenStudy (anonymous):

ok ill post it